A projectile is launched horizontally from a 20-m tall edifice with a vox of 25 m/s. How long will it take for the projectile to hit the ground?


A projectile is launched horizontally from a 20-m tall edifice with a vox of 25 m/s. Calculate the range.


A projectile is launched horizontally from a 20-m tall edifice with a vox of 25 m/s. What is the velocity of the projectile along the x axis right before hitting the ground?


A projectile is launched horizontally from a 20-m tall edifice with a vox of 25 m/s. What is the final velocity of the projectile along the y axis right before hitting the ground?


1
Expert's answer
2021-01-30T06:26:34-0500

a)

"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot20\ m}{9.8\ \dfrac{m}{s^2}}}=2.02\ s."

b)

"x=v_{0x}t=25\ \dfrac{m}{s}\cdot2.02\ s=50.5\ m."

c) The horizontal velocity of the projectile (velocity along "x" axis) remains unchainged during the flight. Thus, "v_x=25\ \dfrac{m}{s}."

d)

"v_y=v_{0y}-gt=0-9.8\ \dfrac{m}{s^2}\cdot2.02\ s=-19.79\ \dfrac{m}{s}."

The sign minus means that the velocity of the projectile directed downward.